OpenStudy (anonymous):

Choose the ordered pair (x,y) that is a solution of the equation. -15x – 2y = 23 A. (-3,11) B. (2,-9) C. (3,-11) D. (-2,9)

OpenStudy (anonymous):

plug the points in the equation the one which will satify the equation is the answer lets try A (-3,11) put x=-3 and y =11 -15(-3)-2(11)=23 45-22=23 23=23 wow we got it right so it must be A